# Unveiling the Capabilities of Large Language Models in Detecting Offensive Language with Annotation Disagreement

### v2

@@ Abstract
Unveiling the Capabilities of Large Language Models in Detecting Offensive Language with Annotation Disagreement
Junyu Lu1, Kai Ma1, Kaichun Wang1, Kelaiti Xiao1 Roy Ka-Wei Lee2, Bo Xu1, Liang Yang1, Hongfei Lin1 1Dalian University of Technology, 2Singapore University of Technology and Design dutljy@mail.dlut.edu.cn, hflin@dlut.edu.cn
@@ 3.1 Evaluation of Binary Classification Performance
We evaluate the accuracy for offensive and non-offensive language across different agreement levels, as shown in Figure 1. We observed that across all agreement levels, LLMs demonstrate higher accuracy in identifying offensive language than non-offensive language. In particular, for low-agreement non-offensive samples ($N^{0}$), accuracy drops to 45.77%, indicating a strong tendency to misclassify ambiguous content as offensive. This over-sensitivity could lead to false positives in automated moderation systems, increasing the risk of justified content removal and restricting legitimate speech.
@@ 3.2 Evaluation of Relationship between Agreement Degree and LLM Confidence
We analyze how well LLM confidence aligns with human annotation agreement, as a well-calibrated model should exhibit high confidence for clear cases and lower confidence for ambiguous cases. If LLMs assign high confidence to disagreement samples, this may indicate overconfidence, limiting their ability to reflect human-like uncertainty. To evaluate this, we apply the self-consistency method Chen and Mueller (2024); Wang et al. (2023b), which resamples model outputs under varying temperature settings to estimate confidence.
@@
We assess self-consistency using Cohen’s Kappa ($\kappa$), measuring how stable LLM outputs remain across multiple sampling attempts. As shown in Figure 2, self-consistency decreases for lower agreement samples but remains above 0.75 even for $A^{0}$ cases, indicating strong internal agreement. While high self-consistency is desirable for clear-cut cases, it becomes problematic in ambiguous cases, as it suggests that LLMs remain overconfident even when human annotators disagree. This rigidity limits the model’s ability to adjust for nuanced linguistic or contextual differences.
@@ 4.2 Impact of Disagreement Samples on Instruction Fine-tuning
We replicate the instruction fine-tuning experiment with Qwen2.5-7B using the same training and test data. The results closely align with those of LLaMa3-7B, confirming that these insights generalize across different model architectures. See Appendix B.4 for details.
@@ 5 Related Work
In recent years, large language models (LLMs) have rapidly emerged, showcasing extensive world knowledge and strong reasoning capabilities Kojima et al. (2022); Ouyang et al. (2022); OpenAI (2023). Many researchers have proposed diverse tasks to deeply analyze the relationship between the model’s outputs and human judgments Xu et al. (2024); Fan et al. (2024). In addition, the confidence of LLMs in their outputs has also attracted attention from researchers, which is often used to assess the reliability and robustness of the generated content Jiang et al. (2021). Various methods for estimating confidence have been proposed Zhang et al. (2020); Wang et al. (2023b); Tian et al. (2023); Lin et al. (2022). In this study, we employ the most straightforward approach, self-consistency, to estimate the model’s confidence.
@@
Researchers have developed various methods for detecting offensive language Founta et al. (2018); Davidson et al. (2017); Mathew et al. (2021). As research advances, many studies argue that treating offensive language detection as a binary classification is an idealized assumption Basile et al. (2021); Basile (2020); Plank (2022), as annotation disagreement are inherent in datasets for such subjective task Pavlick and Kwiatkowski (2019); Uma et al. (2021b). Using majority voting for annotation agreement leads to information loss Davani et al. (2022), as these disagreements arise from the subtlety of the samples, not labeling errors Uma et al. (2022). Leonardelli et al. (2023) emphasizes that detection models should recognize this disagreement, rather than just improving classification performance.
@@ 6 Conclusion
Key findings of this work include: (1) a systematic evaluation of LLMs on annotation disagreement, (2) insights into how disagreement samples improve learning, and (3) guidelines for leveraging disagreement-aware training strategies. These results emphasize the need for model calibration techniques to mitigate overconfidence and for training strategies that incorporate disagreement to improve generalization. Future research should explore dynamic fine-tuning approaches and confidence-aware moderation systems to bridge the gap between LLM decisions and human subjectivity.
Limitations
@@
(3) In evaluating the confidence of LLMs, we adopt a straightforward approach based on temperature resampling. We have noted another common method, the Logit-based approach Guo et al. (2017); Zhang et al. (2020), which involves using the logits of category-specific tokens to compute statistical probabilities within the model’s output. This method may provide deeper insights into the decision-making mechanisms of LLMs when handling disagreement samples. We plan to explore and evaluate this method in future work.
Ethics Statement
The opinions and findings contained in the samples of this paper should not be interpreted as representing the views expressed or implied by the authors. Accessing the MD-Agreement dataset requires users to agree to the creators’ usage agreements. The usage of these samples in this study fully complies with these agreements.
@@ B.2 Consistency Analysis Across Different LLMs
B.5 Error Analysis
@@
To gain deeper insight into the challenge posed by offensive language with annotation disagreement, we manually inspect the set of samples misclassified by the models. The following two main types of errors are identified, with samples and predictions from GPT-4o and LLaMa3-72B shown in Table B3 for illustration:

### v3

# Is LLM an Overconfident Judge? Unveiling the Capabilities of LLMs in Detecting Offensive Language with Annotation Disagreement

@@ Abstract
Is LLM an Overconfident Judge? Unveiling the Capabilities of LLMs in Detecting Offensive Language with Annotation Disagreement
Junyu Lu1, Kai Ma1, Kaichun Wang1, Kelaiti Xiao1,2 Roy Ka-Wei Lee3, Bo Xu1, Liang Yang1, Hongfei Lin1thanks: Corresponding author. 1Key Laboratory of Social Computing and Cognitive Intelligence, Dalian University of Technology 2School of Computer Science and Technology, Xinjiang Normal University 3Social AI Studio, Singapore University of Technology and Design dutljy@mail.dlut.edu.cn, hflin@dlut.edu.cn
@@ 3.1 Evaluation of Binary Classification Performance
We evaluate the accuracy for offensive and non-offensive language across different agreement levels, as shown in Figure B3. We observed that across all agreement levels, LLMs demonstrate higher accuracy in identifying offensive language than non-offensive language. In particular, for low-agreement non-offensive samples ($N^{0}$), accuracy drops to 45.77%, indicating a strong tendency to misclassify ambiguous content as offensive. This over-sensitivity could lead to false positives in automated moderation systems, increasing the risk of justified content removal and restricting legitimate speech.
@@ 3.2 Evaluation of Relationship between Agreement Degree and LLM Confidence
We analyze how well LLM confidence aligns with human annotation agreement, as a well-calibrated model should exhibit high confidence for clear cases and lower confidence for ambiguous cases. If LLMs assign high confidence to disagreement samples, this may indicate overconfidence, limiting their ability to reflect human-like uncertainty. To evaluate this, we apply the self-consistency method Wang et al. (2023b); Chen and Mueller (2024), which resamples model outputs under varying temperature settings to estimate confidence.
@@
We assess self-consistency using Cohen’s Kappa ($\kappa$), measuring how stable LLM outputs remain across multiple sampling attempts. As shown in Figure 2, self-consistency decreases for lower agreement samples but remains above 0.75 even for $A^{0}$ cases, indicating strong internal agreement (see Table A2). While high self-consistency is expected for unambiguous cases, it becomes problematic in ambiguous cases where human annotators disagree. This suggests that LLMs fail to capture the inherent uncertainty, exhibiting overconfidence by treating ambiguous inputs as clear-cut. Such rigidity hinders their ability to adapt to nuanced linguistic and contextual variations.
@@ 4.2 Impact of Disagreement Samples on Instruction Fine-tuning
We replicate the instruction fine-tuning experiment with Qwen2.5-7B using the same training and test data. The results closely align with those of LLaMa3-7B, confirming that these insights generalize across different model architectures (see Appendix B.4). Additionally, we conduct an in-depth analysis of the generalization ability of the trained models on out-of-distribution (OOD) data using disagreement samples, drawing on the bias-variance trade-off theory (see Appendix B.5).
@@ 5 Related Work
In recent years, large language models (LLMs) have rapidly emerged, showcasing extensive world knowledge and strong reasoning capabilities Kojima et al. (2022). Many researchers have proposed diverse tasks to deeply analyze the relationship between the model’s outputs and human judgments Xu et al. (2024); Fan et al. (2024). In addition, the confidence of LLMs in their outputs has also attracted attention from researchers, which is often used to assess the reliability and robustness of the generated content Jiang et al. (2021). Various methods for estimating confidence have been proposed Wang et al. (2023b); Tian et al. (2023); Lin et al. (2022). In this study, we employ the most straightforward approach, self-consistency Wang et al. (2023b), to estimate the model’s confidence.
@@
Researchers have developed various methods for detecting offensive language Founta et al. (2018); Davidson et al. (2017); Mathew et al. (2021). As research advances, many studies argue that treating offensive language detection as a binary classification is an idealized assumption Basile et al. (2021); Basile (2020); Plank (2022), as annotation disagreement is inherent in datasets for such subjective task Pavlick and Kwiatkowski (2019); Uma et al. (2021b). Using majority voting for annotation agreement leads to information loss Davani et al. (2022), as these disagreements arise from the subtlety of the samples, not labeling errors Uma et al. (2022). Leonardelli et al. (2023) emphasizes that detection models should recognize this disagreement, rather than just improving classification performance.
@@ 6 Conclusion
Key findings of this work include: (1) a systematic evaluation of LLMs on annotation disagreement, (2) insights into how disagreement samples improve learning, and (3) guidelines for leveraging disagreement-aware training strategies. These results emphasize the need for model calibration techniques to mitigate overconfidence and for training strategies that incorporate disagreement to improve generalization.
Future research should explore dynamic fine-tuning approaches and confidence-aware moderation systems to bridge the gap between LLM decisions and human subjectivity. We hope that our findings and insights will contribute to the further development of offensive language detection, LLM-as-a-judge, and human-AI collaboration.
Limitations and Ethics Statement
@@
(3) In evaluating the confidence of LLMs, we adopt a straightforward approach based on temperature resampling, which has been widely used in recent LLM literature, especially for evaluating uncertainty in black-box settings, where internal logits or token probabilities may not be accessible. We fully acknowledge the diversity of uncertainty metrics and plan to explore logit-based calibration methods Guo et al. (2017) (e.g., ECE, Brier score, and temperature scaling) and Bayesian approximations in future work to deepen our understanding of LLM uncertainty in handling annotation disagreements.
(4) The opinions and findings contained in the samples of this paper should not be interpreted as representing the views expressed or implied by the authors. Accessing the MD-Agreement dataset requires users to agree to the creators’ usage agreements. The usage of these samples in this study fully complies with these agreements  [^1] The code of this work is available at https://github.com/DUT-lujunyu/Disagreement.
Acknowledgments
This research is supported by the Natural Science Foundation of China (No. 62376051), the Liaoning Provincial Natural Science Foundation Joint Fund Program (2023-MSBA-003), the Fundamental Research Funds for the Central Universities (DUT24LAB123, DUT24MS003), and the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Award ID: MOE-T2EP20222-0010). Any opinions, and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of the National Natural Science Foundation of China and the Ministry of Education, Singapore.
@@ B.2 Consistency Analysis Across Different LLMs
B.5 In-Depth Analysis of the Generalization Ability of Trained Models
In this section, we further explore the generalization ability of the trained models using disagreement samples and conduct an in-depth analysis from a more fundamental perspective of the bias-variance trade-off theory.
In the few-shot experiments (Table 5 and B1), we observed a slight accuracy drop when incorporating all three levels of annotation agreement (LLaMa-3 using $A^{++/+/0}$: 83.51%) compared to using a more balanced combination (LLaMa-3 using $A^{++/0}$: 83.8%). This aligns with the classic trade-off: including overly complex or ambiguous training data may increase variance, leading to overfitting and reduced generalization. Conversely, training only on unanimous samples may lead to underfitting, as the model may not sufficiently learn nuanced decision boundaries.
To further explore this hypothesis, we fine-tuned LLaMA-3 on subsets of MD-Agreement with different agreement levels and evaluated its generalization performance on the Implicit Hate Corpus dataset (IHC) ElSherief et al. (2021), which contains binary-labeled offensive language samples. Since IHC does not include annotator agreement metadata, we focused purely on classification accuracy as an out-of-distribution (OOD) metric. The experimental results are presented in Table B3.
The best OOD performance was achieved when fine-tuning on medium-agreement ($A^{+}$) samples, suggesting that they strike a favorable balance between challenge and clarity. This finding further confirms that annotation disagreement level can act as a proxy for data complexity, and that moderate complexity helps mitigate both bias and variance.
MD (ID) IHC (OOD)
Model Acc.\;$\uparrow$ MSE\;$\downarrow$ Acc.\;$\uparrow$ F1\;$\uparrow$
LLaMa3-8B 70.92 0.1856 58.87 47.97
$\qquad w/\;A^{++}$ 75.79 0.1671 59.19 48.23
$\qquad w/\;A^{++}$ 77.04 0.1552 59.35 48.41
$\qquad w/\;A^{++}$ 73.99 0.1348 59.05 48.12
Table B3: Performance of fine-tuned LLaMa3-8B on the test set of MD-Agreement and IHC.
Category N. O. Num
Swearing 173 415 588
Threatening 27 18 45
Personal Bias 661 402 1,063
Rhetorical Question 57 30 87
Reported Speech 64 27 91
Word Play 24 22 46
Sarcasm 112 50 162
Analogy 41 45 86
False Assertion 18 12 30
Ungrammatical 10 8 18
No context 258 76 334
Not Complete 6 2 8
Noise 12 0 12
Total 1,463 1,107 2,570
Table B4: Statistics of the MD-Agreement-v2 Sandri et al. (2023), where N. and O. represent non-offensive and offensive samples, respectively.
B.6 Impact of Linguistic Features in Disagreement Samples
We evaluate the performance of LLMs on disagreement samples with varying linguistic features in the zero-shot setting. We use MD-Agreement-v2 Sandri et al. (2023) as the benchmark, which is an extended version of MD-Agreement that provides fine-grained annotations of the linguistic features present in each sample, including Swearing, Threatening language, sensitive topics provoking ambiguity (Personal Bias), Word Play, Rhetorical Questions reflecting the publisher’s viewpoint, Reported Speech, Sarcasm, Analogy, False Assertions, or the use of Ungrammatical words or sentence structures. The definitions of features and dataset statistics are provided in Table B4 and B6. Four widely used models are selected for this assessment, including GPT-4, Claude-3.5, Mixtral 8x22B, and LLaMa-3 70B. The F1-score serves as the evaluation metric. Based on the results shown in Figure B3, the following conclusions can be drawn:
LLM exhibits weak performance in detecting samples with complex, context-dependent linguistic features and noise. When processing samples containing expressions of Sarcasm and Rhetorical Questions, LLMs consistently exhibit significantly poor performance, with average F1 scores of 54.6% and 61.6%, respectively, which notably lower the overall dataset performance of 67.7%. In contrast, for Swearing and Analogy samples, which feature explicit token-level markers like swear words or comparative terms, LLMs achieve F1 scores of 83.8% and 71.2%, respectively. This indicates that samples containing complex linguistic features, which require the model to integrate complete and nuanced context for reasoning and judgment, remain a challenge for existing LLMs. Another lower-performing category is the one containing Ungrammatical expressions, with an F1 score of only 57.1%, reflecting the limited ability of LLMs to handle noise, which further reduces their reliability in real-world text moderation tasks.
Performance varies significantly across different LLMs, even when detecting the same linguistic feature. When detecting samples involving Reported Speech that cites others’ viewpoints, the best-performing model, LLaMa3-70B, achieved an F1 score 16.3% higher than the weakest Mixtral-8$\times$22B. In addition, for samples containing Rhetorical Questions, GPT-4 outperformed Claude-3.5 by 10.3%, despite similar overall performance across the dataset. This underscores the substantial performance differences among models when handling offensive language with the same linguistic feature. These differences stem from variations in the models’ training mechanisms and the datasets used for training. In practical applications, selecting the most suitable model for detecting samples with specific linguistic features may yield better results.
Figure B3: Performance of LLMs in detecting disagreement samples with varying linguistic features, where Overall refers to the MD-Agreement-v2.
@@ B.6 Impact of Linguistic Features in Disagreement Samples
Taxonomy Linguistic Definition Example
Swearing The sample includes swearing words or insults (e.g., "prick", "retarded"). 2nd wave about to be a bitch. ($O^{0}$)
Threatening The sample contains violent expressions or linguistic threats. U r going to jail. ($N^{0}$)
Personal Bias The sample includes specific topics that can be interpreted subjectively (e.g., "racist") or expresses specific critical opinions (e.g., "no vax"). #DemocratsAreDestroyingAmerica #Black Lives Matter is a terrorist organization. ($N^{0}$)
Rhetorical Question The sample contains a question asked to highlight a concept, rather than to seek an answer. Why do we treat our prisoners like this? Is it really because we decided once you commit a crime you’re worthless non-human? ($N^{0}$)
Reported Speech The sample reports something someone else stated (e.g., a newspaper headline). White Bystanders With Rifles Stare Down George Floyd Protesters: ’You Ain’t Got No Guns’. ($O^{0}$)
Word Play The sample uses literary devices that alter certain words, including acronyms, alliterations or puns. The only people ripping this country apart are your fellow liberal #DemocRATS and your militant concubines. ($N^{0}$)
Sarcasm The sample aims to express the opposite of its surface meaning humorously or to mock someone/something. Who knew a side effect of COVID would be gross incompetence. ($N^{0}$)
Analogy The sample contains metaphors or analogies. Trump is a walking petri dish. His goal is to
spread the virus to as many people as possible. ($N^{0}$)
False Assertion The publisher expresses something contrary to his thoughts, or something false and exaggerated in relation to the context. Another attempt backfired on them, George Floyd cured Covid-19 and opened up the economy! ($N^{0}$)
Ungrammatical The sample includes non-standard varieties, slang, code-switching, or simple typing errors. Mane it’s hard for some of da blaxk people out dere when dey go into a store they got everybody looking at them "what they doing". ($N^{+}$)
No Context The sample contains ambiguous pronouns or links, lacking sufficient context. Dude this guy is serious? And trump retweeted this?? Please anonymous take them out. ($N^{+}$)
Not Complete The sample does not convey complete information or miss some parts. Wtaf, a farce in three parts: ($O^{0}$)
Noise The sample is clearly not offensive but marked as such (or vice versa). In a singular voice, art across the world. ($N^{+}$)
Table B6: Definitions and examples of various linguistics in the offensive language with annotation disagreement, with definitions from Sandri et al. (2023) and examples from the MD-Agreement-v2 dataset. Three categories - No Context, Not Complete, and Noise - arise from the low quality of the data itself and are excluded from LLM evaluation. In contrast, we retain evaluation for the Ungrammatical category, as it reflects non-standard expressions and typos commonly found in offensive language.
B.7 Error Analysis
To gain deeper insight into the challenge posed by offensive language with disagreement, we manually inspect the set of samples misclassified by the models and conduct an error analysis. The following two main types of errors are identified, with samples and predictions from GPT-4o and LLaMa3-72B shown in Table B5 for illustration: